OpenStudy (jordanloveangel):

The equation y = 1.6x represents the number of laps Henry can swim over time, where y is the number of laps and x is time in minutes. The table shows the number of laps Larry can swim over time. How many laps can each boy complete per minute, and who can swim laps at a faster rate? Henry: laps per minute Larry: laps per minute The boy who swims at a faster rate is . (Type H for Henry or L for Larry) Larry Time (min) 2.5 4 5.5 8 Laps completed 4.5 7.2 9.9 14.4

OpenStudy (jordanloveangel):

@Haichi

OpenStudy (perl):

can you make an equation of a line using the data for Larry

OpenStudy (jordanloveangel):

idk how 2 solve the problem

OpenStudy (perl):

first find the slope , pick two points from Larry swim times

OpenStudy (perl):

The ordered pair (x,y) represents ( time , # laps) So we can use points (2.5 , 4.5) and ( 4, 7.2)

OpenStudy (perl):

slope = (difference in y) / (difference in x) So here, slope = (7.2 - 4.5 ) / ( 4 - 2 )

OpenStudy (jordanloveangel):

@TheEdwardsFamily

OpenStudy (theedwardsfamily):

multiply 1.6 times 1.8 to get harry think

OpenStudy (theedwardsfamily):

no I take that back

OpenStudy (theedwardsfamily):

@DanJS

OpenStudy (theedwardsfamily):

@Nnesha @MARC_

OpenStudy (danjs):

?

OpenStudy (theedwardsfamily):

do you have any guess what the answer might be @DanJS ?

OpenStudy (danjs):

@perl did exactly right y = mx + b ; m = slope = laps / min Larry -- y = 1.6x + 0 Henry--- Find slope m = laps / min and compare to larry's laps/min

OpenStudy (danjs):

Which part is unclear?

OpenStudy (jordanloveangel):

im so confused

OpenStudy (danjs):

They both swim at a constant rate. For Larry: Laps = 1.6*(minutes) In 1 minute, how many laps does Larry swim? Let minutes = 1

OpenStudy (theedwardsfamily):

gtgt bed @jordanloveangel! :(

OpenStudy (jordanloveangel):

kk

OpenStudy (jordanloveangel):

Henry: 1.6 laps per minute, Larry: 1.8 laps per minute and the boy who swims at a faster rate is L.